20 tablets weighing 2g was powdered. Exactly 50mg was transferred to 100mL volumetric flask and diluted with water. 10mL aliquot was transferred to another 100mL and diluted to volume. The A value of the solution was 0.652 and the absorptivity value was 15mg/mL/cm. What is the mg of paracetamol per tablet from the following data? Please provide a step by step solution. Thank you.


1
Expert's answer
2021-02-18T12:49:14-0500

Absorption formulae:

A = a b c,

Whereby

A – absorbance (0.652)

a – absorptivity (15mg/ml/cm)

b – length of part (1cm)

c – concentration

substitution;

c = A/(ab) = 0.652/ (15mg/ml/cm x 1cm) = 0.0435mg/ml

Result;

The second 100ml volumetric flask contained 0.0435mg/ml x 100ml = 4.35mg of paracetamol.

Hence the first 100ml flask contained 4.35mg x 10 = 43.5mg of paracetamol. This means that, 50mg of tablet consists of 43.5mg of paracetamol.

Since 20 tablets weigh 2g, one tablet weighs 2g/20 =0.1g = 100mg.

Thus as 50mg of tablet contains 43.5mg of paracetamol, one tablet (100mg) contains 43.5mg x 2 =87 mg of paracetamol.

 

Hence the answer is 87mg